Pakistan cement industry has shown very fast progress in last few years and has
become a leading sector of the economy. At the time of inception, there were only
four cement plants in Pakistan which has now grown to over 30 units. Cement
manufacturers have also expanded their production capacity to meet higher demand
because of construction projects and CPEC. Maple Leaf Cement Factory Limited is
a reputable largest manufacturer of cement in Pakistan. The company was set up in
1956. Suppose the quantity demanded and quantity supplied functions of cement
industry in starting week of current month are:
Qd = 5000-6P
Qs= 1500+P
Where ‘P’ is the price in rupees per bag of cement and ‘Qd’ is quantity demanded of
cement in number of bags. ‘Qs’ is quantity supplied of cement in number of bags.
Requirements:
a. Calculate the market equilibrium level of price and quantity.
1
Expert's answer
2019-05-14T09:07:05-0400

a. The market equilibrium level of price and quantity are:

Qd = Qs,

5000 - 6P = 1500 + P,

7P = 3500,

P = 500,

Q = 1500 + 500 = 2000 units.